What capacitance when connected in series with a 500Ω resistor will limit the current

drawn from a 48-mV 465-kHz source to 20μA?



1
Expert's answer
2021-01-19T12:00:35-0500

Answer

For RC circuit resistance is given by

"Z=\frac{V}{I}=\sqrt{R^2+X_c^2}"

"X_c=\frac{1}{2\pi fC}" -------eq1

So using above equation capacitance can be written as

"X_c=\sqrt{\frac{V^2}{I^2}-R^2}"

"X_c=551\times10^4 ohm" -------eq2

Using equation 1 and 2

Capacitance

"C=\frac{1}{2\times3.14\times465\times10^3\times23.47\times10^2}"

C=0.15nF